year 1974 nominal gdp is 2000, gdp deflator is 100, year 1975 nominal gdp is 3000 and gdp deflator is 120, year 1976 nominal gdp is 3750 and gdp deflator is 150, year 1977 nominal gdp is 6,000 and gdp deflator is 200. This economy experienced inflation from 1974 to 1975, from 1975 to 1976 to 1977.
Answer #1

Ans:

Inflation from 1974 to 1975 is 20%

Explanation:

= { ( gdp deflator in 1975 - gdp deflator in 1974 ) / gdp deflator in 1974 } * 100

= {( 120 - 100 ) / 100 }*100

= ( 20 / 100 ) * 100

= 0.2 * 100 = 20%

Inflation from 1975 to 1976 is 25%

Explanation:

= { ( gdp deflator in 1976 - gdp deflator in1975 ) / gdp deflator in 1975 } * 100

= {( 150 - 120 ) / 120 }*100

= ( 30 / 120 ) * 100

= 0.25 * 100 = 25%

Inflation from 1976 to 1977 is 33.33%

Explanation:

= { ( gdp deflator in 1977 - gdp deflator in1976 ) / gdp deflator in 1976 } * 100

= {( 200 - 150 ) / 150 }*100

= ( 50 / 150 ) * 100

= 0.3333 * 100 = 33.33%